Grade 12Physical Chemistry

Energy needed to ionise a H atom if electron occupies n=5 orbit?

Aoishi
9 Years agoGrade 12

1 Answer

DEEPENDRA SINGH
9 Years ago
given-atomic number of H (z)=1
n=5
so,from the formula of energy we get-
13.6*z2/n2=13.6*1/25=0.544 J